OpenStudy (anonymous):

Find the radius of the circle whose diameter has the given endpoints. Then use your answers to find the standard form of the equation of the circle. The endpoints of the diameter are (2,7) and (6,9). The center is (4,8). The radius is ?. The equation in standard form is ?

OpenStudy (kira_yamato):

rad = sqrt 5 units Eqn: \[\left[ x-4 \right]^2 + \left[ y-8 \right]^2 = 5\]

OpenStudy (anonymous):

Thanks